OpenStudy (anonymous):

A bottle of liquid medication contains 300mL and costs $10.00. If the usual dose is 20mL, how much does each dose cost?

OpenStudy (saifoo.khan):

300 : 10 20 : x cross multiply.

OpenStudy (anonymous):

(20/300) * 10 = $0.67

OpenStudy (saifoo.khan):

0.67

OpenStudy (anonymous):

300/10 = 20/x x=$0.67